OpenStudy (anonymous):

Mary has $3.00 in nickels, dimes, and quarters. If she has twice as many dimes as quarters and twenty-seven more nickels than dimes, how many coins of each type does she have?

OpenStudy (anonymous):

x+y+z=3 y=2z x=y+27 solve this

OpenStudy (anonymous):

and it comes out to 3 seperate answers?

OpenStudy (anonymous):

i need to know how many quarters, dimes, and nickels she has left of each

OpenStudy (anonymous):

how many of each type she has*

OpenStudy (anonymous):

here d+n+q=3 2d=q 27+d=n now you can solve, by substituting

OpenStudy (anonymous):

plug in the value for q, and n in terms of d into the d+n+q equation, you will get the number of d/dimes. than plug that number in to the two other equations to get the number of q/quarters and n/nickels

OpenStudy (anonymous):

what does it come out to?

OpenStudy (anonymous):

i cant figure out how to do it based on what you are saying

OpenStudy (anonymous):

ok just tell me what q is someone. i will solve the rest i just need one variable if no one can let me check them and answer by answering them all

OpenStudy (anonymous):

ok fixed this\[0.10d + 0.05n +0.25q = 3.00\] 2d=q 27+n = d \[0.10d + 0.05(27+d) +0.25(2d) = 3.00\] \[(0.10+0.05+0.50)d +1.35=3.00\]

OpenStudy (anonymous):

\[(0.10+0.05+0.50)d = 1.65\]

OpenStudy (anonymous):

solve for d, sorry i don't have a calculator this will be easier for you to solve

OpenStudy (anonymous):

once you have d, plug that value into the two short equations, 2d=q and 27+d = n

OpenStudy (anonymous):

so d = .0025 based on whats in the parnetheses. where is the 1.65 from

OpenStudy (anonymous):

d should equal 1.65/(0.10+0.05+0.50)

OpenStudy (anonymous):

so would be closer to a value of 3, but I may be solving this inaccurately

OpenStudy (anonymous):

1.65 from 3.00-1.35

OpenStudy (anonymous):

1.65 * .0025 = .004125?

OpenStudy (anonymous):

sorry the first equation is: 5x+10y+25z=300 y=2z x=y+27 solution: x=33 y=6 z=3 33 nickels, 6 dimes and 3 quarters

OpenStudy (anonymous):

THANKYOU

OpenStudy (anonymous):

easier to use whole numbers instead of decimals, but you should be able to get same answers

OpenStudy (anonymous):

i don't live in U. S. so i didn't know the value in cents of nickels, dimes and quarters